OpenStudy (anonymous):

What are the factors of 25x2 + 40x + 16? (5x + 4)(5x + 4) (5x − 4)(5x − 4) (5x + 2)(5x + 8) (5x − 8)(5x − 2)

OpenStudy (anonymous):

(5x + 4)(5x + 4)

OpenStudy (anonymous):

please explain

OpenStudy (anonymous):

Use the quadratic formula (-b +- sqrt (b^2 - 4ac)) / (2a) where the equation is in the form of ax^2 + bx + c = 0

OpenStudy (anonymous):

25x^2+40x+16=0

OpenStudy (anonymous):

Or you can try this 25x^2 + 40x + 16 = 25x^2 + 20x + 20x + 16 = 5x(5x + 4) + 4(5x + 4) = (5x + 4)(5x + 4)

OpenStudy (anonymous):

well thank you ! :)